OpenStudy (anonymous):

What is the surface area of a rectangular prism with length fifteen millimeters, width eleven millimeters , and height three millimeters? fifty-eight square millimeters one hundred seventy-four square millimeters four hundred eighty-six square millimeters four hundred ninety-five square millimeters

OpenStudy (anonymous):

@calculusxy

OpenStudy (calculusxy):

2(15*11) 2(11*3) 2(15*3) Find the products, then add.

OpenStudy (calculusxy):

Surface area is really the total amount of area that each of the faces have .

OpenStudy (anonymous):

dont get it

OpenStudy (calculusxy):

Okay did you get the products?

OpenStudy (anonymous):

245,025

OpenStudy (calculusxy):

No I meant from the multiplication equations that I gave you.

OpenStudy (anonymous):

yeah i know

OpenStudy (calculusxy):

2(15*11)= 2(11*3)= 2(15*3)= What are the products for these?

OpenStudy (calculusxy):

Did u multiply them by two?

OpenStudy (anonymous):

330 66 90

OpenStudy (calculusxy):

So now add.

OpenStudy (anonymous):

add the numbers i got

OpenStudy (calculusxy):

What did u get?

OpenStudy (anonymous):

486

OpenStudy (anonymous):

@calculusxy